OpenStudy (anonymous):

A triangle has side lengths 10, 15, and 7. Is the triangle acute, obtuse, or right? Show your work and explain.

OpenStudy (campbell_st):

use pythagoras' theorem right angle \[c^2 = a^2 + b^2\] obtuse is \[c^2 > a^2 + b^2\] and acute if \[c^2 < a^2 + b^2\] you have c = 15, a = 10 and b = 7 hope it helps

OpenStudy (jdoe0001):

did you get what campbell_st said?